# The Preliminary Mauve Science Programme: Science themes identified for the first year of operations–ACRONYMS The Preliminary Mauve Science Programme: Science themes identified for the first year of operations

### v2

@@ 4.5.2 Variability of Herbig Ae/Be stars and implications for accretion and planet formation
Monitoring the spectrophotometric modulation of light curves provides a powerful diagnostic of accretion in young stars, as variability often reflects changes in the accretion geometry or rate. Variability studies are therefore an excellent tool for probing accretion physics across stellar masses. For CTTSs, it is well established that accretion from the circumstellar disk to the stellar surface is regulated by a strong magnetic field (Fig. 9), which truncates the disk at approximately five stellar radii above the surface (Calvet and Gullbring, 1998). Material follows magnetic field lines onto the star, impacting in compact regions–accretion “footpoints”–that cover a small fraction of the stellar surface. If the stellar dipole is misaligned with the rotation axis, or if higher-order field components dominate (Gregory and Donati, 2011), the footpoints become localised, producing rotationally modulated photometric variability as they move in and out of view (Rebull et al., 2020). However, if the magnetic field is weak or absent, as is thought to be the case for most Herbig Ae/Be stars, the situation changes drastically. Extrapolating the field geometry inward suggests that accretion could occur over a much larger area of the stellar surface, potentially forming an equatorial accretion ring in the limit of negligible magnetic confinement. In this scenario, the accretion emission would show weak or no rotational modulation, an important observational signature that Mauve is well equipped to detect. The continuum emission from accretion shocks in CTTSs typically has a temperature of around 10,000 K (Valenti et al., 1993), similar to the effective temperature of an A0 star. At first glance, this similarity might make accretion emission in Herbig Ae/Be stars difficult to detect. However, Fairlamb et al. (2015) demonstrated that the accretion excess remains clearly visible in the spectral region shortward of the Balmer jump, providing a reliable diagnostic for accretion even in these hotter stars (Fig. 10). Mauve is particularly well suited to study this phenomenon across a large number of targets. Simulations using MauveSim (Saba et al., 2025) for the Herbig Ae star AB Aur (V = 7.05) show that in a 300-second exposure, the accretion excess shortward of the Balmer jump is readily detectable. In Fig. 11 the expected photospheric emission of the A0 star (shown in orange in the simulation) contrasts clearly with the purple excess due to accretion. By monitoring the variability of this excess emission over time, we can directly probe in the accretion rate and geometry in Herbig Ae/Be stars. In addition to the blue continuum, accretion also produces strong hydrogen recombination emission lines, particularly H and other Balmer lines, as gas flows along magnetic channels or accretes directly from the disk. Although Mauve’s low spectral resolution will broaden these lines as shown in Fig. 12, flux variations in H remain a sensitive and independent tracer of accretion variability. Combining line and continuum diagnostics will provide a comprehensive picture of the accretion processes in these intermediate-mass stars.
@@ 4.6 Binaries in Exotic Stellar Populations
From the observational perspective, high-quality photometric and spectroscopic data are crucial for accurately determining the properties of exotic stars and discerning their formation mechanisms. Therefore, ultraviolet imaging and spectroscopy offer a unique window to identify hot companions to the above-mentioned stars if present, thereby helping to constrain their formation and differences, if any, between the field and clusters populations. The near-ultraviolet spectra will complement the existing spectra in visible to estimate stellar parameters (, , ) using multi-wavelength spectral energy distributions, test evolutionary models, and derive the properties of binary components if present. The other advantage of observing in ultraviolet is the large amplitude variations compared to optical and infrared regimes in the case of pulsating stars, and stars exhibiting any kind of stellar activity. ultraviolet spectra will also enable the discernment of the evolutionary pathways of hot stars in the field and compare their properties with those in star clusters.
@@
Any ultraviolet excess which might indicate the presence of a hot companion.
@@
Flux variability will be studied to determine if there are Keplerian motions (eclipses).
@@
Activity indicators will be used to disentangle stellar activity from Keplerian motions.
@@ 5 Discussion & Conclusions
In this paper we have presented the science themes prioritised by the Mauve Science Team for the first year of Mauve’s operations. Initial simulations with MauveSim suggest strong performance across the different scientific goals and observational approaches presented here. However, the actual performance of Mauve will be confirmed through in-orbit testing and calibration, therefore, as the mission evolves, Mauve’s scientific focus and observational strategy will need to remain flexible. New members are expected to join the survey collaboration throughout Year 1, with additional observational hours released periodically. Future science themes may focus on the detailed characterisation of accretion signatures in young, low-mass stars, as illustrated by the simulation in Fig. 15. Future publications will expand on Mauve’s science operations, data analysis, satellite performances and science results.
@@ Acknowledgements
T. A. A. Sigut is grateful for support from the Fredrick Hunt Physics and Astronomy Fund and the Natural Sciences and Engineering Research Council of Canada (NSERC) through the Discovery Grant program. C. Dong acknowledges the funding support from NASA grant 80NSSC23K1115 and the Alfred P. Sloan Research Fellowship. E. Pancino and S. Rani are co-funded by the European Union (ERC-2022-AdG, "StarDance: the non-canonical evolution of stars in clusters", Grant Agreement 101093572, PI: E. Pancino). K. Vida’s research is funded by the Hungarian National Research, Development and Innovation Office grant Élvonal KKP-143986.

### v3

@@ 4.5.2 Variability of Herbig Ae/Be stars and implications for accretion and planet formation
Monitoring the spectrophotometric modulation of light curves provides a powerful diagnostic of accretion in young stars, as variability often reflects changes in the accretion geometry or rate. Variability studies are therefore an excellent tool for probing accretion physics across stellar masses. For CTTSs, it is well established that accretion from the circumstellar disk to the stellar surface is regulated by a strong magnetic field (Fig. 10), which truncates the disk at approximately five stellar radii above the surface (Calvet and Gullbring, 1998). Material follows magnetic field lines onto the star, impacting in compact regions–accretion “footpoints”–that cover a small fraction of the stellar surface. If the stellar dipole is misaligned with the rotation axis, or if higher-order field components dominate (Gregory and Donati, 2011), the footpoints become localised, producing rotationally modulated photometric variability as they move in and out of view (Rebull et al., 2020). However, if the magnetic field is weak or absent, as is thought to be the case for most Herbig Ae/Be stars, the situation changes drastically. Extrapolating the field geometry inward suggests that accretion could occur over a much larger area of the stellar surface, potentially forming an equatorial accretion ring in the limit of negligible magnetic confinement. In this scenario, the accretion emission would show weak or no rotational modulation, an important observational signature that Mauve is well equipped to detect. The continuum emission from accretion shocks in CTTSs typically has a temperature of around 10,000 K (Valenti et al., 1993), similar to the effective temperature of an A0 star. At first glance, this similarity might make accretion emission in Herbig Ae/Be stars difficult to detect. However, Fairlamb et al. (2015) demonstrated that the accretion excess remains clearly visible in the spectral region shortward of the Balmer jump, providing a reliable diagnostic for accretion even in these hotter stars (Fig. 11). Mauve is particularly well suited to study this phenomenon across a large number of targets. Simulations using MauveSim (Saba et al., 2025a) for the Herbig Ae star AB Aur (V = 7.05) show that in a 300-second exposure, the accretion excess shortward of the Balmer jump is readily detectable. In Fig. 12 the expected photospheric emission of the A0 star (shown in orange in the simulation) contrasts clearly with the purple excess due to accretion. By monitoring the variability of this excess emission over time, we can directly probe in the accretion rate and geometry in Herbig Ae/Be stars. In addition to the blue continuum, accretion also produces strong hydrogen recombination emission lines, particularly H and other Balmer lines, as gas flows along magnetic channels or accretes directly from the disk. Although Mauve’s low spectral resolution will broaden these lines as shown in Fig. 13, flux variations in H remain a sensitive and independent tracer of accretion variability. Combining line and continuum diagnostics will provide a comprehensive picture of the accretion processes in these intermediate-mass stars.
@@ 4.6 Binaries in Exotic Stellar Populations
From the observational perspective, high-quality photometric and spectroscopic data are crucial for accurately determining the properties of exotic stars and discerning their formation mechanisms. Therefore, ultraviolet imaging and spectroscopy offer a unique window to identify hot companions to the above-mentioned stars if present, thereby helping to constrain their formation and differences, if any, between the field and clusters populations. The near-ultraviolet spectra will complement the existing spectra in visible wavelengths to estimate stellar parameters (, , ) using multi-wavelength spectral energy distributions, test evolutionary models, and derive the properties of binary components if present. The other advantage of observing in ultraviolet is the large amplitude variations compared to optical and infrared regimes in the case of pulsating stars, and stars exhibiting any kind of stellar activity. ultraviolet spectra will also enable the discernment of the evolutionary pathways of hot stars in the field and compare their properties with those in star clusters.
@@
Any ultraviolet excess which might indicate the presence of a hot companion. Our candidates are mostly low-mass (<2 Msun) stars in various evolutionary phases, like main sequence, subgiant branch, red giant branch, hot subdwarfs. The hot companions are expected to be mostly white dwarfs or extremely low mass stars, with masses mostly in the range 0.1 to 0.5 Msun, thus we will be finding mostly unequal mass binaries.
@@
Flux variability will be studied to determine if there are Keplerian motions (eclipses). Eclipses will be easier to find for close binaries which, depending on the radius of the brighter star, will be on orbits that can range from one to several hundred days. However, shorter periods below about 10 days will be disproportionally more likely to produce eclipses.
@@
Activity indicators will be used to disentangle stellar activity from Keplerian motions. Activity can be a natural result of the stellar properties, such as for example for low mass main sequence stars, but can also be stimulated by close binary interactions. To distinguish between the two, we will use the whole target list, including low-mass stars from other themes in a collaborative effort, to establish the baseline of "normal" activity for different types of stars.
@@ 5 Discussion & Conclusions
In this paper we have presented the science themes prioritised by the Mauve Science Team for the first year of Mauve’s operations. Initial simulations with MauveSim suggest strong performance across the different scientific goals and observational approaches presented here. However, the actual performance of Mauve will be confirmed through in-orbit testing and calibration, therefore, as the mission evolves, Mauve’s scientific focus and observational strategy will need to remain flexible. New members are expected to join the survey collaboration throughout Year 1, with additional observational hours released periodically. Future science themes may focus on the detailed characterisation of accretion signatures in young, low-mass stars, as illustrated by the simulation in Fig. 16. Future publications will expand on Mauve’s science operations, data analysis, satellite performances and science results, including reporting on payload and detector ageing throughout the mission lifetime.
@@ Acknowledgements
The Mauve project has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101082738. T. A. A. Sigut is grateful for support from the Fredrick Hunt Physics and Astronomy Fund and the Natural Sciences and Engineering Research Council of Canada (NSERC) through the Discovery Grant program. C. Dong acknowledges support from NASA grant No. 80NSSC23K1115, the Alfred P. Sloan Research Fellowship, and the IBM Einstein Fellow Fund at the Institute for Advanced Study, Princeton. E. Pancino and S. Rani are co-funded by the European Union (ERC-2022-AdG, "StarDance: the non-canonical evolution of stars in clusters", Grant Agreement 101093572, PI: E. Pancino). K. Vida’s research is funded by the Hungarian National Research, Development and Innovation Office grant Élvonal KKP-143986.
Data Availability
All data for this paper are contained within the article; BSSL is able to share simulated data upon request.